### v4

Abstract
We estimate the numbers of always takers, compliers, defiers, and never takers in the sample of people in an experiment, rather than a hypothetical population from which it was drawn, using structure from the randomization design. Our data include only a binary intervention and outcome. We develop a visualization to show that samples with defiers can sometimes generate the data in more ways than samples without defiers, yielding a higher design-based likelihood. We propose a maximum likelihood decision rule that can harness this evidence, which is not captured by standard hypothesis tests, and we provide optimality conditions. We illustrate the output of our decision rule for all possible data in samples of 50 and 200 with half in intervention, demonstrating a pattern in when the MLE includes defiers despite a positive average effect. We provide insights into effect heterogeneity in two published experiments with interventions that could plausibly backfire for some people despite statistically significant positive average effects on takeup of desirable health behaviors. In both, our 95% credible sets include the estimated Fréchet bounds, demonstrating that evidence is weak. Yet, our MLE includes no defiers in one; in the other, the MLE includes a count of defiers equal to the estimated upper Fréchet bound, over 18% of the sample. The MLE can support a monotonicity assumption or a specific alternative as a step toward improving the average effect of future interventions by targeting them away from noncompliers. Our dbmle package, compatible with Python and Stata, implements our statistics.
Keywords: health, econometrics, statistical decision theory, experiments
JEL Codes: I1, C1, C44, C9
1 Introduction
We begin with a stylized example that demonstrates how a design-based model of an experiment can reveal evidence beyond the average effect. In Section 2, we formalize the model and present a likelihood function for the joint distribution of potential outcomes in the sample, which we use to construct a maximum likelihood decision rule. In Section 3, we apply our maximum likelihood decision rule to two real-world experiments. Section 4 discusses our contributions to the literature. Section 5 presents implications for econometric and applied research.
1.1 Visualization of a Stylized Example: A Design-Based Model Can Reveal Evidence Beyond the Average Effect
Suppose you are considering an intervention to increase takeup of a desired health behavior. For example, the desired health behavior could be to quit smoking, and the intervention could be a payment. You are worried that your intervention could backfire, even if it increases takeup on average. Some people who would otherwise quit smoking could be dissuaded from doing so by a payment. You are also concerned that you would rather not waste payments on people who would quit smoking or continue smoking regardless.
A colleague tells you that you can use an experiment to estimate the average effect of the intervention in the population, but it cannot provide evidence that the intervention backfires if the average effect is positive. You are worried that even if the estimated average effect is positive, the intervention could backfire for some people and be ineffectual for others. Rather than focusing on the average effect in the population, you focus your attention on the distribution of effects within the fixed sample of subjects in an experiment.
You design an experiment with six people such that you randomly assign exactly three people to intervention and three people to control. You run the experiment, and you estimate the marginal distributions of outcomes in intervention and control: two people have taken up the behavior in intervention, and one person has taken up the behavior in control. These results seem promising because takeup is higher in intervention than control, yielding an average effect of 1/3 on a control takeup rate of 1/3.
Why did you observe the data you observed? What would you have seen under the other possible randomized assignments? Can the experiment reveal any evidence about effect heterogeneity? In more technical terms, what is the joint distribution of “potential outcomes” (neyman1923; rubin1974; rubin1977) in the sample? Counterfactual questions like these have attracted recent interest in the study of causal inference (gelman2013; pearl2018; imbens2020; dawid2022).
We assume, as in design-based inference, that the potential outcomes of each person are fixed, and therefore the joint distribution of potential outcomes in the sample is fixed, before randomization. Following angrist1996, we classify people into types based on their potential outcomes in intervention and control: “always takers” take up the desired health behavior in intervention and control (effect=0); “compliers” take up in intervention but not control (effect=1); “defiers” do not take up in intervention but take up in control (effect=-1); and “never takers” do not take up in intervention or control (effect=0). We specify the joint distribution of potential outcomes with the numbers of always takers, compliers, defiers, and never takers in the sample. The joint distribution of potential outcomes provides more information than the average effect; it provides the full distribution of effects and disambiguates between zero effects from always and never takers.
What is your best estimate of the numbers of always takers, compliers, defiers, and never takers in the experiment you ran with six people? We estimate with our design-based maximum likelihood decision rule that the sample includes four compliers and two defiers, and we develop a novel visualization of potential outcomes to illustrate the logic behind our answer in Figure 1. We represent each of the six people in the sample with a circle. The left half of each circle represents the potential outcome in intervention, and the right half represents the potential outcome in control. Orange represents “takeup,” white represents “no takeup,” and grey represents “unobserved.”
Rectangles with lines down the middle represent experiments with intervention on the left and control on the right. Inside each experiment, assignment to intervention reveals the potential outcome in intervention—the left of the rectangle reveals the left of the circle, which can be orange for takeup or white for no takeup—but the potential outcome in control is unobserved, so it is grey. Similarly, assignment to control reveals the potential outcome in control—the right of the rectangle reveals the right of the circle, which can be orange for takeup or white for no takeup—but the potential outcome in intervention is unobserved, so it is grey.
The data from your experiment, depicted in the box at the top of Figure 1, show takeup among two of three people in intervention and one of three people in control. Therefore, the estimated marginal distribution of the potential outcome in intervention is 2/3 takeup, and the estimated marginal distribution of the potential outcome in control is 1/3 takeup. These two marginal distributions define the “estimated Fréchet set.” We refer to all joint distributions of potential outcomes with the same marginal distributions as members of the same “Fréchet set.”
The three rows of Figure 1 give the three joint distributions of potential outcomes with the same marginal distributions as the estimated Fréchet set: half circles in which the potential outcomes are unobserved in the data are filled in so that they have the same marginal distributions as the data. Four of the six circles have orange on the left (takeup occurs among 2/3 of people in intervention), and two of the six circles have orange on the right (takeup occurs among 1/3 of people in control). As shown in the three rows, there are three ways to combine the colored half circles into colored full circles, each yielding a different number of defiers (circles with white on the left and orange on the right). This depiction demonstrates how the estimated Fréchet set determines the estimated Fréchet bounds on defiers, which extend here from a lower bound of 0 to an upper bound of 2. How do you decide among these three possible estimates?
Figure 1: In an Experiment with Six People, the Likelihood Varies Among Joint Distributions of Potential Outcomes with the Same Marginal Distributions as the Data (2/3 Takeup in Intervention, 1/3 Takeup in Control) and is Maximized with Four Compliers and Two Defiers
The joint distribution of potential outcomes in the first row is consistent with the imbens1994 assumption of “monotonicity” in the sample, whereby there can be compliers or defiers, but not both. Since the estimated effect of 1/3 is positive, you assume away defiers. You then deduce that the person who takes up in control must be an always taker and the person who does not take up in intervention must be a never taker. Assuming that the randomization guarantees that the count of each type is the same in each arm, the sample includes two always takers and two never takers. The remaining two people must be compliers, so the joint distribution of potential outcomes consistent with the monotonicity assumption includes two always takers, two compliers, and two never takers. However, you are uncomfortable assuming monotonicity because you are concerned that there could be defiers, and you would prefer to decide based on evidence.
The joint distributions of potential outcomes in the next two rows include defiers. The one in the second row includes one always taker, three compliers, one defier, and one never taker. The one in the third row includes four compliers and two defiers. There is a rationale in the literature for deciding on the distribution in the second row—it has one defier, the midpoint (li2019) of the Fréchet bounds on defiers. It can therefore seem surprising that we decide on the distribution in the third row with two defiers.
Each row in Figure 1 provides a rationale for our decision using a “design-based model” that combines the design of the randomization with a model of potential outcomes to specify a data generating process and use it to generate all possible data. As depicted in the top of each row, there are 20 possible randomized assignments of six people that result in three in intervention ($6\text{ choose }3=20$). Each of the six people in the experiment has a separate row. The person’s type and randomized assignment to intervention or control determines the data the experiment reveals, which can be counterfactual to what you observed. The assignments that generate the data you observed—2/3 takeup in intervention and 1/3 takeup in control—are in lighter shading.
The number of randomized assignments that generate the data you observed gives the numerator of the design-based likelihood, reported in the third column. Paraphrasing the baby board book “Statistical Physics for Babies” (ferrie2017), which provides style inspiration for our visualization, physicists refer to the number of ways that you could have seen what you have seen as “entropy.” The design-based likelihood is equal to the entropy—the number of assignments that generate the data (shown in lighter shading)—divided by the number of possible assignments. The first row, with zero defiers, generates the data in eight of 20 possible assignments, so the likelihood is $8/20=40\%$. The second row, with one defier, generates the data in six of 20 possible assignments, so the likelihood is $6/20=30\%$. The last row, with two defiers, generates the data in 12 of 20 possible assignments, so the likelihood is $12/20=60\%$.
Why do some joint distributions of potential outcomes generate the data in more randomized assignments than others and therefore have higher likelihoods than others? The bottom half of each row sorts the six people by type and separately shows how the always takers, compliers, defiers, and never takers must be randomized into intervention and control to generate the data. In the first row, there are three types—always takers, compliers, and never takers—and there must be a balance between intervention and control within each type to generate the data. In the second row, there are four types—always takers, compliers, defiers, and never takers—and there must be an imbalance between intervention and control within each type to generate the data. In the third row, there are only two types—compliers and defiers—and there must be a balance between intervention and control within each type to generate the data.
Likelihoods are higher when the people who generate the data are as similar as possible in two senses: they belong to fewer types, and they have a counterpart of the same type in the opposite arm such that those of the same type are balanced between intervention and control. The people who generate the data are most similar in the third row: they belong to the fewest number of types (only two), and those of the same type are balanced between intervention and control. Comparison of the first and third rows illustrates how having fewer types can increase the likelihood. In the first row, there are two ways to randomize one of two always takers into intervention (2 $2\text{ choose }1$=2), two ways to randomize one of two compliers into intervention (2 $2\text{ choose }1$=2), and two ways to randomize one of two defiers into intervention (2 $2\text{ choose }1$=2), so there are eight assignments that generate the data ($2\times 2\times 2=8$). In the third row, there are six ways to randomize two of four compliers into intervention ($4\text{ choose }2=6$) and two ways to randomize one of two defiers into intervention (2 $2\text{ choose }1$=2), so there are 12 assignments that generate the data ($6\times 2=12$). Joint distributions of potential outcomes other than those in Figure 1 can generate the data (the true joint distribution of potential outcomes need not preserve the estimated marginal distributions), but by exhaustive grid search, you determine that the global maximizer of the likelihood is the one depicted in the third row, which only includes two types.
By our proposed maximum likelihood decision rule, you conclude that your best estimate for the joint distribution of potential outcomes in the sample consists of four compliers and two defiers. Using this distribution, you recover the estimated average effect of 1/3—of the six people in the experiment, the four compliers have an effect of positive one, and the two defiers have an effect of negative one ($(4\times 1+2\times-1)/6=1/3$). The design-based model of the experiment enables you to learn about not just the average effect of your intervention, but also about something more primitive: the joint distribution of potential outcomes, which yields the entire distribution of effects.
2 A Design-Based Model of an Experiment
2.1 A Design-Based Model Implies a Likelihood Function for the Joint Distribution of Potential Outcomes in the Sample
In this section, we formalize the design-based model of an experiment presented in our stylized example. A sample consists of $n$ subjects, each of whom has a binary potential outcome in intervention $Y_{I}\in\{0,1\}$ and a binary potential outcome in control $Y_{C}\in\{0,1\}$, where $1$ represents takeup and 0 represents no takeup. A subject’s observed outcome depends only on their own potential outcomes and their inclusion in the intervention or control arm, ruling out network-type effects through a “no interference” (cox1958) or “stable unit treatment value” (rubin1980) assumption.
Subjects are randomly assigned to intervention ($Z=I$) or control ($Z=C$), and assignment to intervention $Z$ is independent of a subject’s potential outcomes $(Y_{I},Y_{C})$. The observed outcome $Y$ is:
$Y=\mathbf{1}_{\{Z=I\}}(Y_{I})+\mathbf{1}_{\{Z=C\}}(Y_{C}),$ (1)
where $\mathbf{1}_{\{\cdot\}}$ is the indicator function.
Subjects in the sample belong to one of four types or “principal strata,” defined by combinations of potential outcomes (frangakis2002). Let $\theta_{11}$ represent the number of always takers $(Y_{I}=1,Y_{C}=1)$, $\theta_{10}$ the number of compliers $(Y_{I}=1,Y_{C}=0)$, $\theta_{01}$ the number of defiers $(Y_{I}=0,Y_{C}=1)$, and $\theta_{00}$ the number of never takers $(Y_{I}=0,Y_{C}=0)$. The collection of these four counts ${\boldsymbol{\theta}}=(\theta_{11},\theta_{10},\theta_{01},\theta_{00})$ constitutes the joint distribution of potential outcomes in the sample, which we write as counts rather than shares to ease the following likelihood notation. The experimental data ${\boldsymbol{X}}=(X_{I1},X_{I0},X_{C1},X_{C0})$ consist of the counts of subjects who take up in intervention $X_{I1}$, who do not take up in intervention $X_{I0}$, who take up in control $X_{C1}$, and who do not take up in control $X_{C0}$.
We adopt a “design-based” approach, in which the joint distribution of potential outcomes $\boldsymbol{\theta}$} is fixed, but unknown, and all randomness in the experimental data $\boldsymbol{X}$} comes from the random assignment of subjects to either intervention or control. As shown in Appendix A, randomization implies a distribution for the experimental data $\boldsymbol{X}$} given the joint distribution of potential outcomes $\boldsymbol{\theta}$}, yielding a likelihood expression. For Bernoulli randomization, also known as simple randomization or independent and identically distributed (IID) randomization, with probability $p$ of assignment to intervention, the likelihood is:
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $=\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{i}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-i}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-i}$
$\qquad\qquad\times\binom{\theta_{00}}{x_{I0}+x_{C1}+i-\theta_{11}-\theta_{01}}$
$\qquad\qquad\times p^{x_{I1}+x_{I0}}(1-p)^{x_{C1}+x_{C0}},$
where the set $\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})$ restricts $i$ such that the binomial coefficients remain well defined.
For a completely randomized experiment with $m$ subjects randomized into intervention, the likelihood is:
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $=\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{i}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-i}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-i}$
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}\bigg/\binom{n}{m}$.
While these randomization procedures imply different distributions for the experimental data, once we condition on the experimental data, the likelihoods are both proportional to:
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $\propto\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{i}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-i}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-i}$
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}$. (2)
This last expression appears in copas1973.
This likelihood function coincides with the likelihoods we derive in the introductory example of Section 1. For the first distribution with zero defiers and the third distribution with two defiers in Figure 1, there is a single term in the likelihood summation. For the second distribution with one defier, there are two.
2.2 The Design-Based Likelihood Can Vary with the Number of Defiers within the Fréchet Bounds Determined by the Estimated Marginal Distributions of Potential Outcomes
In the previous section, we present a likelihood for the joint distribution of potential outcomes in the sample. Each joint distribution of potential outcomes implies marginal distributions of potential outcomes in the intervention and control arms. The marginal distribution of the potential outcome in intervention is the number of subjects for whom $Y=1$ in intervention:
$\theta_{1\bullet}\equiv\theta_{11}+\theta_{10}.$
The marginal distribution of the potential outcome in control is the number of subjects for whom $Y=1$ in control:
$\theta_{\bullet 1}\equiv\theta_{11}+\theta_{01}.$
Given a pair of marginal distributions of potential outcomes in intervention and control, there are numerous joint distributions consistent with them. boole1854, hoeffding1940, and frechet1957 derive bounds on the possible copulas connecting marginal distributions of random variables into joint distributions. By applying these bounds to a pair of marginal distributions of potential outcomes, we can determine the set of all joint distributions of potential outcomes consistent with the pair of marginal distributions. We refer to the set of joint distributions of potential outcomes consistent with a given pair of marginal distributions as the “Fréchet set” of those marginal distributions, $\mathcal{F}(\theta_{1\bullet},\theta_{\bullet 1};n)$. In Figure 1, the three rows comprise the Fréchet set for $\theta_{1\bullet}=4$ and $\theta_{\bullet 1}=2$. We index the distributions in a Fréchet set by their number of defiers, $\theta_{01}$, which can take values in the range
$\max\{0,-(\theta_{1\bullet}-\theta_{\bullet 1})\}\leq\theta_{01}\leq\min\{\theta_{\bullet 1},n-\theta_{1\bullet}\}.$ (3)
These inequalities provide the so-called “Fréchet bounds” on the number of defiers. Within these bounds, each value for the number of defiers implies a unique number of always takers, never takers, and compliers consistent with the given marginals:
$\mathcal{F}(\theta_{1\bullet},\theta_{\bullet 1};n)=\Big\{{\boldsymbol{\theta}}\in\mathbb{Z}^{4}_{+}:\,$ $\theta_{01}\in[\max\{0,-(\theta_{1\bullet}-\theta_{\bullet 1})\},\min\{\theta_{\bullet 1},n-\theta_{1\bullet}\}],$
$\theta_{10}=\theta_{1\bullet}-\theta_{\bullet 1}+\theta_{01},$
$\theta_{11}=\theta_{\bullet 1}-\theta_{01},$
$\theta_{00}=n-\theta_{1\bullet}-\theta_{01}\Big\}.$
The bounds on the number of defiers $\theta_{01}$ ensure that these counts are non-negative and sum to $n$.
We can form estimates of the marginal distributions of potential outcomes in intervention and control directly using the observed takeup rates in intervention and control:
$\hat{\theta}_{1\bullet}({\boldsymbol{x}})$ $=n\left(\frac{x_{I1}}{x_{I1}+x_{I0}}\right)$
$\hat{\theta}_{\bullet 1}({\boldsymbol{x}})$ $=n\left(\frac{x_{C1}}{x_{C1}+x_{C0}}\right).$ (4)
When we report these estimates, we round them if necessary for comparison to our count-based distributions. In our design-based model of an experiment, these are unbiased estimators of the sample marginal distributions of potential outcomes for a completely randomized experiment (before rounding). In a standard sampling-based model of an experiment, dividing by $n$ yields consistent estimates of the population marginal distributions of potential outcomes.
It is well-known that, in the sampling-based model of an experiment, the takeup counts in intervention and control contain no information about the joint distribution of potential outcomes in the population beyond their marginal distributions. In Appendix B, we recreate this result by showing that the likelihood of the population-level joint distribution of potential outcomes is constant, conditional on the population-level marginal distributions. In other words, the likelihood function is flat across every population-level Fréchet set.
In our design-based model of an experiment, however, the data do contain information about the sample joint distribution of potential outcomes beyond the sample marginal distributions. That is, the likelihood function varies with the number of defiers within sample-level Fréchet sets. Figure 2 plots the likelihood values of each joint distribution of potential outcomes in the Fréchet set formed by the estimated marginal distributions from the second empirical example that we will discuss in Section 3. We index the joint distributions in this Fréchet set by their number of defiers along the horizontal axis. The height of the bar is the likelihood of the joint distribution in the Fréchet set with the specified number of defiers. Within this Fréchet set, the likelihood is highest at the distribution with 21 defiers, the upper Fréchet bound on defiers. As the number of defiers increases, the likelihood initially decreases, and then increases. In all empirical examples we have considered, the likelihood maintains a U-shape within the estimated Fréchet set and all other Fréchet sets, and is maximal at either the minimum or maximum number of defiers. If we normalize likelihoods in this Fréchet set to sum to one, the smallest collection of likelihoods to contain 95% of the mass excludes 9 defiers, indicated by the lighter shading.
Figure 2: The Likelihood Varies with the Number of Defiers within the Estimated Fréchet Set in Johnson and Goldstein (2003), And The Smallest Collection of Likelihoods to Contain 95% of the Mass Excludes 9 Defiers (in Lighter Shading)
2.3 The Design-Based Likelihood Enables Our Proposed Maximum Likelihood Decision Rule
We consider a statistical decision problem in the style of wald1949, whereby the decision maker wishes to correctly guess the joint distribution of potential outcomes in the sample, having viewed the outcome of the experiment. In Appendix C.1, we present the formal decision problem. Statistical decision theory provides two benefits in our setting: first, decision theory is natural to apply in our design-based setting, unlike alternative criteria like consistency that depend on large-sample or asymptotic assumptions; second, statistical decision theory provides straightforward methods to quantify the gains from exploiting the variation in our likelihood over other decision rules that ignore such variation.
We propose choosing the joint distribution of potential outcomes that maximizes the design-based likelihood. Let $\widehat{{\boldsymbol{\Theta}}}({\boldsymbol{x}})$ be the set of joint distributions $\boldsymbol{\theta}$} that maximize the likelihood given the observed data $\boldsymbol{x}$}:
$\widehat{{\boldsymbol{\Theta}}}({\boldsymbol{x}})=\arg\max_{{\boldsymbol{\theta}}}\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$
There are finitely many vectors of four integers that sum to the number of participants in the experiment $n$, so $\widehat{{\boldsymbol{\Theta}}}({\boldsymbol{x}})$ is nonempty. We define the maximum likelihood decision rule as an estimator $\widehat{{\boldsymbol{\theta}}}_{MLE}$ that places equal weight on each maximizer of the likelihood:
$\mathbb{P}(\widehat{{\boldsymbol{\theta}}}_{MLE}({\boldsymbol{x}})={\boldsymbol{\theta}}\mid{\boldsymbol{x}})=\begin{cases}\dfrac{1}{\#\{\widehat{{\boldsymbol{\Theta}}}({\boldsymbol{x}})\}},&\text{if}\quad{\boldsymbol{\theta}}\in\widehat{{\boldsymbol{\Theta}}}({\boldsymbol{x}}),\\ 0,&\text{otherwise,}\end{cases}$
where $\{\bullet\}$ is the counting measure. When the likelihood is unimodal, the decision rule chooses the maximizer with probability one. We solve this optimization problem through an exhaustive grid search over the finitely many joint distributions of potential outcomes consisting of $n$ subjects.
2.4 Justifications for our Proposed Maximum Likelihood Decision Rule
Our primary motivation for proposing the maximum likelihood decision rule is that maximum likelihood rules are considered “standard” in econometric settings, and maximum likelihood estimators are familiar to practitioners. In the following subsections, we provide three additional justifications: first, that the maximum likelihood decision rule varies systematically with the data; second, that the decision rule is Bayes optimal under the appropriate conditions; and third, that the decision rule maximizes entropy, and therefore “assumes the least” (jaynes1968) about the joint distribution of potential outcomes.
2.4.1 The Maximum Likelihood Decision Rule Varies Systematically with the Data
Direct inspection of the likelihood in (2) does not reveal obvious patterns in the maximum likelihood estimate. However, by computing the maximum likelihood estimate for every possible realization of the data in fixed sample sizes, we can observe clear patterns in the estimates. Figure 3 illustrates how the maximum likelihood estimates from our proposed decision rule vary systematically with all possible realizations of the data in completely randomized experiments of 50 and 200 subjects with half in intervention. In a sample of 200 with half in intervention, there are $10,201=101\times 101$ possible realizations of the data and $1,373,301={203\choose 3}$ possible joint distributions of potential outcomes, so the visualization reflects results of grid searches over $14,009,043,501=10,201\times 1,373,301$ total elements.
In the subfigure for each sample size, the horizontal axis gives the takeup count in intervention, and the vertical axis gives the takeup count in control. Each cell represents a possible realization of the data, which specifies an estimated Fréchet set in terms of takeup count in intervention and control. The visualization demonstrates that an estimated Fréchet set can also be specified with the estimated average effect, which increases diagonally downward, and the total takeup count in the sample, which increases diagonally upward. An estimated Fréchet set can also be specified with an estimated average effect and a takeup rate in control, as is common in empirical practice.
The figure provides a visualization of our statistical decision rule by reporting statistics from our maximum likelihood estimates of the numbers of always takers, compliers, defiers, and never takers in the sample. Maximum likelihood estimates in cells within dotted outlines include the same types: always takers (A), compliers (C), defiers (D), and never takers (N). Darker purple shading within cells indicates higher counts of defiers in the maximum likelihood estimates.
Figure 3: Visualization of Our Proposed Statistical Decision Rule: How the MLE Varies with All Possible Data from Samples of 50 and 200 with Half in Intervention
The figure shows clear patterns in how the MLEs vary with the data, which are visibly similar when we increase the sample size by a factor of four. MLEs in the corners only include one type: always takers if takeup is full in intervention and control, compliers if takeup is full in intervention and zero in control, defiers if takeup is zero in intervention and full in control, and never takers if takeup is zero in intervention and control. MLEs along the edges and diagonals only include the two types in the corners they connect.
In the convex lens-shaped region around the upward-sloping diagonal, MLEs only include always and never takers: the two types in the null hypothesis of the fisher1935 exact test. In samples of 50 and 200, the region in which Fisher’s exact test fails to reject the null hypothesis at the 5% level is depicted within the grey dashed line. It is larger than the region in which the MLE includes only always and never takers in both sample sizes.
The figure illustrates how our design-based maximum likelihood decision rule reveals evidence that is obscured by Fisher’s exact test. Fisher’s exact test only considers the null hypothesis that the joint distribution of potential outcomes includes only always and never takers. If it rejects, it does not suggest an alternative. If it fails to reject, a joint distribution of potential outcomes that includes types other than always and never takers can maximize the likelihood. With our use of statistical decision theory, we can harness the richer evidence conveyed in the figure beyond the grey dashed lines.
In the convex lens-shaped region around the downward sloping diagonal, MLEs only include compliers and defiers. This region is related to the region in which an exact test of the null hypothesis that everyone is affected does not reject. There are some realizations of the data near the center of the region in which Fisher’s exact test fails to reject the null hypothesis that no one is affected, but our maximum likelihood estimates indicate that everyone is affected. These estimates are relevant to empirical practice for researchers who wonder if they obtained a near-zero estimated average effect not because no one was affected but because everyone was affected.
Sometimes, the MLE can be a tie between more than one distribution. MLEs are unique in the sample of 50 with half in intervention, but two-way ties occur in the sample of 200. In cells bisected by the lines indicated with asterisks, the midlines in which takeup is exactly half in intervention or control, ties occur between MLEs in cells on either side of the line. Only at the edges of the figure do the ties result in a unique MLE because the MLE is the same on either side of the line. For experiments with intervention and control arms unequal in size, ties appear in less predictable patterns.
The MLEs include three types in all other regions, with two regions for each combination. Even though the maximizer of the likelihood only includes two types in our stylized example from Figure 1, maximization of the likelihood sometimes requires trading off between the higher likelihood of fewer types and the higher likelihood of more balance between intervention and control within each type. If restricting the sample to two types requires too much imbalance between intervention and control, the maximizer of the likelihood includes three types. [^1] Viewing Pascal’s triangle as a representation of $N$ choose $K$ can help us to visualize the tradeoff between fewer types (moving to the next row of the triangle, increasing $N$ within each type) and more balance within each type (moving to the middle of the triangle within a row, moving $K$ closer to $N/2$ within each type). We thank Elizabeth Ananat for this point. There are no regions in which a distribution that maximizes the likelihood contains all four types. Therefore, many joint distributions of potential outcomes never maximize the likelihood for any realization of the data. For example, if a sample of 50 consisted of 18 always takers, 22 compliers, 8 defiers, and 2 never takers, the most probable realization of the data would be takeup of 20 in intervention and 13 in control, but the corresponding MLE estimate of 26 always takers, 14, compliers, 0 defiers, and 10 never takers would miss the 8 defiers entirely. We have observed MLEs that contain all four types in experiments with intervention and control arms unequal in size, although we have only observed them tied with MLEs that contain three types.
The region outlined in bold, in which the estimated average effect is positive, is of particular interest because MLEs in the shaded portion of the region include both defiers and compliers, contrary to monotonicity assumptions, as in the stylized example in Figure 1. In all even-sized samples up to 200, the MLE includes defiers when takeup is below half in control and above half in intervention, unless takeup is zero in control or full in intervention. The experiment depicted in Figure 1 falls into the region with defiers because the takeup rate is 1/3 in control and 2/3 in intervention. As shown, the MLE always includes defiers when the estimated average effect is greater than half, unless takeup is zero in control or full in intervention, and it can include defiers even when Fisher’s exact test rejects at the 5% level.
The shading demonstrates that the MLE reveals evidence beyond the average effect: holding the estimated average effect constant along upward sloping diagonal lines, MLEs can include zero or many defiers. copas1973 claims that the likelihood is always maximized at a joint distribution of potential outcomes that preserves the direct estimates of the marginal distributions of potential outcomes from equation (4) and which has either the maximal or minimal number of defiers. In practice, we often find that the likelihood is maximized at a distribution that deviates slightly from these marginal estimates.
2.4.2 The Maximum Likelihood Decision Rule is Bayes Optimal
Our proposed maximum likelihood decision rule is Bayes optimal under appropriate specifications of the decision maker’s utility function and prior. For the utility function, we specify a discrete utility function that assigns zero utility for incorrect guesses and an arbitrary amount of strictly positive utility for correct guesses; for the prior, we specify a prior that is uniform over all possible distributions of potential outcomes for $n$ subjects. Appendix C.2 reviews the formal definition and proof of Bayes optimality under these conditions. [^2] We thank Andriy Norets and Thomas Wiemann for bringing the standard conditions that establish Bayes optimality to our attention.
As we review in Appendix C.2, Bayes optimal decision rules can be derived for alternative utility functions and priors through the numerical solution of a linear programming problem, but we offer motivations for the conditions we have chosen, acknowledging their subjectivity. For the utility function, one motivation is that we can interpret it as the payoff function in a game that rewards “being correct,” as in the well-known Monty Hall game, where the contestant receives a prize only if they guess the correct door. Under this utility function, the maximum likelihood decision rule is the Maximum A Posteriori (MAP) rule, which chooses the mode of the posterior distribution. Under a uniform prior, the posterior distribution is proportional to the likelihood, and the MAP estimate equals the MLE estimate.
The desirable feature of our uniform prior is that it assigns equal weight to all distributions in the same Fréchet set, so it is “uninformative” in the sense that it does not confer any information about the joint distribution of potential outcomes beyond their marginal distributions. Furthermore, we can view our uniform prior as the result of sampling from an underlying population over which we hold an uninformative hyperprior. In sampling settings, it is common to assume a Dirichlet hyperprior over population shares. We choose a Dirichlet hyperprior over the population shares of always takers, compliers, defiers, and never takers with all concentration parameters set to one, which is equivalent to an uninformative uniform distribution for the population. Under standard assumptions on sampling from the population, the potential outcomes of subjects are independently and identically distributed (IID), and the sample distribution of potential outcomes follows a multinomial law conditional on the population distribution. Combining the Dirichlet hyperprior and the sampling process yields a Dirichlet-multinomial prior over the counts of each type in the sample. The Dirichlet-multinomial prior with concentration parameters set to one is equivalent to our uniform prior. In our stylized example in Figure 1, our uniform prior implies that the prior is the same in all three rows, so the updated posterior is proportional to the likelihood.
Regardless of the choice of prior, the ratio of the prior beliefs between any two distributions updates as long as the likelihood ratio differs from one, including within Fréchet sets. For example, consider an alternative prior that is uniform over each type for each person in the sample. This prior arises from a highly informative degenerate hyperprior that assumes the population shares of each type are equal. The resulting prior over sample distributions varies within the sample Fréchet sets, and is therefore also informative. In our stylized example in Figure 1, the resulting prior is 0.0220 for the first row, 0.0293 for the second row, and 0.0037 for the third row. The updated posterior distribution maintains curvature within the Fréchet set and updates towards the distribution in the third row that contains two defiers. To achieve a flat posterior distribution across the rows depicted in Figure 1, the prior for the first row with no defiers would have to be exactly 0.75 (=(20/8)/(20/6)) times as large as the second row with one defier and 1.5 (=(20/8)/(20/12)) times as large as the prior for the third row with two defiers.
A Bayesian perspective allows us to construct credible sets for the joint distribution of potential outcomes. In Appendix D, we describe our algorithm for constructing smallest credible sets for the joint distribution of potential outcomes. Smallest credible sets are difficult to characterize because they are multidimensional, so we also describe an algorithm for constructing smallest credible sets for the number of each type individually.
Bayes optimality ensures that our proposed decision rule does at least as well as all other decision rules under these optimality conditions. To quantify the gains from utilizing the optimal rule, we compare the Bayes expected utility obtained under the optimal rule to two alternative decision rules. The first alternative decision rule, which we refer to as the “Fréchet rule,” estimates the marginal distributions of potential outcomes directly according to (4), and then chooses randomly among the joint distributions of potential outcomes in the corresponding Fréchet set. The Fréchet rule $\widehat{{\boldsymbol{\theta}}}_{F}$ satisfies:
$\mathbb{P}(\widehat{{\boldsymbol{\theta}}}_{F}({\boldsymbol{x}})={\boldsymbol{\theta}})=\begin{cases}\dfrac{1}{\#\mathcal{F}(\hat{\theta}_{1\bullet}({\boldsymbol{x}}),\hat{\theta}_{\bullet 1}({\boldsymbol{x}});n)}&\text{if}\;{\boldsymbol{\theta}}\in\mathcal{F}(\hat{\theta}_{1\bullet}({\boldsymbol{x}}),\hat{\theta}_{\bullet 1}({\boldsymbol{x}});n),\\ 0&\text{otherwise}.\end{cases}$
This decision rule captures the behavior of a statistician who assumes that the experimental data are uninformative within Fréchet sets and is therefore indifferent between any two distributions within the same Fréchet set.
The second alternative decision rule, which we refer to as the “monotonicity rule,” approximates the behavior of a statistician who uses the monotonicity assumption of imbens1994 in the population to estimate the share of compliers or defiers in the sample. The rule uses the observed difference in takeup rate between the intervention and control arms $\Delta$ to approximate the share of compliers or defiers. In particular, if the difference is non-negative, the rule imposes a no-defiers monotonicity restriction, and if the difference is non-positive, it imposes a no-compliers monotonicity restriction.
Under this decision rule, the inferred number of compliers or defiers is given by
$\hat{\theta}_{10}=n\cdot\max\{\Delta,0\},\quad\hat{\theta}_{01}=n\cdot\max\{-\Delta,0\}$
rounded to the nearest whole number. Note that when intervention and control arms are the same size, these expressions are always whole numbers. The remaining types, always takers and never takers, are inferred by assuming balance of type shares across both arms. In the case where there are compliers and no defiers, every treated individual in control must be an always taker, and because assignment is randomized, the proportion of always takers in control should be equal to the proportion of always takers in the sample, giving the estimate $\hat{\theta}_{11}=nx_{C1}/(x_{C1}+x_{C0})$. The remaining individuals must all be never takers. Conversely, if there are defiers and no compliers, every treated individual in intervention must be an always taker, so the estimate of always takers is $\hat{\theta}_{11}=nx_{I1}/(x_{I1}+x_{I0})$ and the remaining individuals are never takers. In summary, the monotonicity rule $\widehat{{\boldsymbol{\theta}}}_{M}$ chooses
$\widehat{{\boldsymbol{\theta}}}_{M}({\boldsymbol{x}})$ $=\big(\hat{\theta}_{11},\hat{\theta}_{10},\hat{\theta}_{01},\hat{\theta}_{00}\big)$
$=\bigg(\textbf{1}_{\{\Delta>0\}}\left[n\frac{x_{C1}}{x_{C1}+x_{C0}}\right]+\textbf{1}_{\{\Delta<0\}}\left[n\frac{x_{I1}}{x_{I1}+x_{I0}}\right],\ n\cdot\max\{\Delta,0\},\ n\cdot\max\{-\Delta,0\},$
$\qquad n-\hat{\theta}_{11}-\hat{\theta}_{01}-\hat{\theta}_{10}\bigg)$
with probability one.
We calculate the exact Bayes expected utility achieved by our optimal maximum likelihood rule and each of the two alternative rules. For a given sample size, we enumerate each of the finitely many feasible sample distributions of potential outcomes $\boldsymbol{\theta}$}. For each distribution $\boldsymbol{\theta}$}, we compute the expected utility of our decision rules by comparing the distribution of estimates they produce under each data realization $\boldsymbol{x}$} to the true distribution. We average the resulting utilities over the values of $\boldsymbol{x}$} according to the likelihood. We then average these expected utilities according to a uniform prior for $\boldsymbol{\theta}$} to get the Bayes expected utility (see Appendix C for additional detail).
Figure 4 plots the ratio of the Bayes expected utility achieved by our proposed maximum likelihood rule to the Bayes expected utility achieved by each of the two alternative rules for even sample sizes between 2 and 200. The gains from using the optimal rule over either of the alternatives increase with the sample size. In a sample size of 200, the Bayes expected utility of the proposed maximum likelihood rule is 1.79 times higher than that of the Fréchet rule, and 1.56 times higher than that of the monotonicity rule.
Figure 4: Under Optimality Conditions, For Increasing Sample Sizes, Performance of Proposed Maximum Likelihood Rule Increases Relative to Fréchet and Monotonicity Rules
2.4.3 The Maximum Likelihood Decision Rule Maximizes Entropy
We can also motivate our maximum likelihood rule through the principle of maximum entropy (jaynes1957a; jaynes1957b), which unites the theory of information with statistical physics. In Section 1.1, we discuss how differences in entropy drive variation in the likelihood such that the maximum likelihood distribution is also the maximum entropy distribution. jaynes1968 advocates for application of the principle of maximum entropy to circumvent the subjective specification of a prior. To make Bayesian decision-making more objective, one option is to choose the least informative prior. Jaynes’ alternative to make the process more objective is to choose the least informative updated distribution. The distribution that can generate the data in the greatest number of ways—the distribution that maximizes entropy, also the distribution that maximizes our design-based likelihood—is the least informative updated distribution. In the Monty Hall game, a celebrated application of Bayesian decision making, it is possible to make the same decision without specifying a Bayesian prior using a maximum likelihood decision rule and the principle of maximum entropy (wang2016).
3 Applications: Counting Defiers in Health Care
We apply our rule to two published experiments with positive, statistically significant average effects on takeup of desired health behaviors and plausible defiers. Figure 3 summarizes the results from all possible empirical applications in experiments with given sizes, illustrating empirical conditions under which our rule yields defiers. It is predictable from the data for each application and the patterns in Figure 3 that our first application yields no defiers and our second application yields defiers, which we confirm with exhaustive grid search. Our focus here is to provide examples to illustrate how applied researchers can present and interpret quantitative results and to provide substantive insights. The top of Table 1 summarizes the context and design for both applications. The statistics in the table can be calculated using our Python- and Stata-compatible dbmle package, documented at https://pypi.org/project/dbmle/ (christy2025dbmle).
In tappin2015, the sample includes 612 pregnant smokers, and the definition of “takeup” is to quit smoking, which is a desired health behavior given concerns that smoking can cause pregnancy complications and lasting adverse effects on children. The experiment uses a completely randomized design, with half of the subjects assigned to an intervention arm with a payment and the other half assigned to a control arm with usual care. Monotonicity assumptions have a long history in the context of interventions that induce pregnant women to quit smoking. In a study discussed by angrist1996, permutt1989 proposed an early monotonicity assumption requiring that women who would quit smoking in control would also quit smoking in an arm that received a multi-faceted intervention. However, literature in psychology and economics expresses concern that payments can backfire in many contexts, including on-time pickup of children from day care (gneezy2000), blood donation (mellstrom2008; lacetera2010), and vaccination (schneider2023). Defiers who would quit smoking in control but not in intervention are plausible in tappin2015 because a payment could backfire by crowding out the intrinsic motivations of pregnant women to quit smoking based on moral ideals and health goals.
The average effect of the payment relative to usual care is a statistically significant 14 percentage point increase in the smoking quit rate on a base of 8% in control. Our proposed decision rule reveals evidence beyond the average effect: our maximum likelihood estimate includes $52/612=8\%$ always takers (effect=0), $86/612=14\%$ compliers (effect=1), and $474/612=77\%$ never takers (effect=0). The weighted average effect across the three types ($8\times 0+14\times 1+77\times 0$) is equal to the estimated average effect of 14 percentage points, which is also equal to the share of compliers. We would obtain the same results if we assumed monotonicity in the sample, so our rule provides evidence in favor of monotonicity without assuming it.
We emphasize that even though our proposed decision rule provides evidence that there are no defiers in the sample, there is considerable uncertainty in our maximum likelihood estimate. As discussed in Appendix D, we report a 95% smallest credible set on the number of defiers in the sample, which we construct by reporting the collection of numbers of defiers occurring within the 95% smallest credible set on the joint distribution of potential outcomes in the sample. The smallest credible set on defiers includes up to 71 defiers, which represents 71/612=12% of the sample size.
Table 1: Standard Statistics and Proposed Design-Based Maximum Likelihood Estimates
Building on the evidence provided by our decision rule, researchers could go on to make the stronger assumption of monotonicity and proceed as usual. With additional data on covariates, they could calculate average characteristics of always takers, compliers, and never takers (imbens1997; katz2001; abadie2002; abadie2003). With additional data on a second stage outcome like birth weight, they could interpret the resulting instrumental variable estimate as a local average treatment effect on compliers (imbens1994).
Our second example, johnson2003, is a high-profile experiment in which a simple change from an opt-in default to an opt-out default increases organ donation almost twofold from 43% in control to 82% in intervention. Results from johnson2003 have inspired state policy changes to encourage life-saving organ donations to combat chronic organ shortages (kessler2025). However, in a recent experiment by kessler2025, in which takeup represents actually becoming an organ donor instead of hypothetically becoming an organ donor, the average effect is close to zero. Especially given the hypothetical outcome in johnson2003, both compliers and defiers are plausible. Compliers may think of the default as an endorsement by the experimenter and thus follow it regardless of what it is. Defiers may think of a default as a threat to their autonomy and thus go against it regardless of what it is.
Given the plausibility of defiers in this context, a monotonicity assumption might not be reasonable, but there is no obvious alternative, so standard practice is to engage with bounds on defiers, which can be large. As shown at the bottom of Table E.1, in which we report auxiliary statistics, the smallest possible number of defiers in the sample is 0 because it is always possible for all people who take up to be always takers and all people who do not take up to be never takers, as in the null of Fisher’s exact test. The largest possible number of defiers in the sample occurs if all 11(=6150)11(=61-50) people who do not take up in intervention and all 23 people who take up in control are defiers, yielding a maximum defier share of $30\%(=(23+11)/115$). The estimated Fréchet bounds are tighter because they incorporate the estimated marginal distributions of potential outcomes in each arm using the observed takeup rates. To calculate the upper Fréchet bound from (3), we consider the bounds on the share of defiers implied by the takeup rates in intervention and control. The intervention takeup rate implies that the share of defiers in intervention and thus in the sample is at most $18\%(=100\%-82\%)$, and the control takeup rate implies that the share of defiers in control and thus in the sample is at most 43%, so the share in intervention binds, and the upper Fréchet bound is 18%.
The Fréchet bounds are estimated with error, so there is even greater uncertainty in the number of defiers. The maximum number of defiers in our 95% smallest credible set is 34, which represents 30% of the sample of 115. The 95% confidence interval on the population fraction of defiers from imbens2004, incorporating the adjustment recommended by jun2023 with pretesting at the 0.01 level, also accounts for estimation of the Fréchet bounds, so we report it for comparison. It is slightly narrower, extending from 0% to 27%, but still includes the estimated Fréchet bounds. Accounting for estimation of the Fréchet bounds is empirically important. As demonstrated in Figure 2, if we do not account for estimation of the Fréchet bounds and instead construct a 95% smallest credible set on defiers within the estimated Fréchet set, variation in the likelihood allows us to exclude 9 defiers— a count in the middle of the Fréchet bounds translated into whole numbers from 0 to 21.
The maximum likelihood estimate indicates that the sample includes $21/115=18\%$ defiers, which is equal to the estimated upper Fréchet bound. The maximum likelihood estimate also includes $28/115=24\%$ always takers, $66/115=57\%$ compliers, and no never takers. The maximum likelihood estimate is the same as the estimate that would be obtained under an assumption of no never takers. Under an assumption of no defiers, the estimated average effect of 39% is entirely due to compliers, who represent 39% of the sample. The maximum likelihood estimate preserves the estimated average effect ($39\%=57\%\times(1)+18\%\times(-1)$), but it tells a very different story. Compliers representing 57% of the sample are affected in one direction and defiers representing 18% of the sample are affected in the other. The total share of people affected in the maximum likelihood estimate of $75\%=(57\%+18\%)$ is almost two times as large as the number of people affected under monotonicity.
In this application, evidence from the design-based model suggests a weaker alternative assumption to monotonicity—an assumption of no never takers—that can subsequently be imposed. Monotonicity can, of course, still be imposed with acknowledgment that it is a stronger assumption. While the assumption of no never takers is still strong, by the principle of maximum entropy, it “assumes the least” (jaynes1968) about the joint distribution of potential outcomes. From the design-based likelihood in (2), the assumption of no never takers can generate the data in $(28\text{ choose }15)\times(66\text{ choose }35)\times(21\text{ choose }11)=8.5\times 10^{31}$ randomized assignments representing 3.4% of the total randomized assignments ($(115\text{ choose }61)=2.5\times 10^{33}$). The assumption of no defiers can generate the data in only $(49\text{ choose }26)\times(45\text{ choose }24)\times(21\text{ choose }11)=7.8\times 10^{31}$ randomized assignments, representing 3.1% of the total randomized assignments. An ex post justification for the assumption of no never takers is that though experimental subjects might feel comfortable complying with or defying defaults imposed by the experimenter, they might feel uncomfortable revealing to the experimenter that they would not donate their organs under either default, especially since they do not know when they are asked a hypothetical question if a later question will ask them for a response under the other default. If an assumption of no never takers is not palatable, evidence for defiers could be used to motivate methods for examining effect heterogeneity, like the collection of additional data or the use of machine learning methods.
If a second stage outcome such as actual organ donation is available, under an assumption of no never takers, the instrumental variable estimate is no longer interpretable as an average treatment effect on compliers without further assumptions. As discussed by angrist1996, under an additional assumption that average treatment effects are the same for compliers and defiers, the instrumental variable estimate is again interpretable as a LATE. dechaisemartin2017 and tchetgen2024 propose alternative assumptions to interpret the instrumental variable estimand in the presence of compliers and defiers.
The value of an assumption of no never takers becomes apparent if data are available on covariates. Under the assumption of no never takers, it is possible to label some specific people as compliers and others as defiers: all people in intervention who do not take up must be compliers, and all people who do not take up in control must be defiers. This reasoning allows us to make claims that the intervention had a causal effect on some specific people in the sample. Using terminology from pearl1999, under an assumption of no never takers, the intervention was a “necessary” cause of no takeup for all 11 people assigned intervention who do not take up, and was a “sufficient” cause of no takeup for all 31 people assigned control who do not take up. Under a monotonicity assumption, it is not possible to directly label any specific people as compliers or defiers because people who take up in intervention can be compliers or always takers, and people who do not take up in control can be compliers and never takers. An assumption that allows us to label some specific people as compliers and others as defiers could be an important first step to learn their characteristics, improve targeting in future samples, and obtain a larger average effect.
4 Contributions to the Literature
Our primary insight is that a randomized experiment can reveal evidence on an estimand more primitive than the average effect: the joint distribution of potential outcomes in the sample. The evidence can support a monotonicity assumption or inform a specific alternative assumption. We obtain evidence using the structure of the implemented randomization process to construct a design-based causal model and its resulting likelihood function, and we evaluate this evidence in a statistical decision theory framework. Our work contributes to the literature on design-based econometrics, statistical decision theory, and causal inference without monotonicity.
In his 1935 book on the design of experiments, Fisher notes that the design of Charles Darwin’s experiments was “greatly superior” to the “statistical methods available at the time” (fisher1935). In their 2017 handbook chapter on the econometrics of experiments, Susan Athey and Guido Imbens write that they “recommend using statistical methods that are directly justified by randomization, in contrast to the more traditional sampling-based approach that is commonly used in econometrics” (athey2017). The sampling-based approach assumes that people in an experiment are randomly sampled from a population, and it attempts to learn about the population. The design-based approach, which we adopt, attempts to learn about the sample of people in the experiment, which is useful since participation in experiments and clinical trials can be far from random (alsan2025) and some samples do not come from a meaningful population (abadie2020).
Our use of a design-based model rather than a sampling-based model enables our contributions. Our ability to learn about defiers with a design-based likelihood may seem paradoxical because it is well known that the traditional sampling-based likelihood does not vary with the number of defiers within the Fréchet bounds determined by the estimated marginal distributions of potential outcomes in intervention and control. As we review in Appendix B, for a given pair of marginal distributions, the likelihood function in the traditional sampling-based model is flat over all copulas connecting them. A large literature has focused on specifying what we can learn from the nonparametric Fréchet bounds in the traditional sampling-based model (balke1997; heckman1997; manski1997mixing; tian2000; zhang2003; imbens2004; fan2010; mullahy2018; li2019; bai2024; semenova2024).
The design-based model allows us to construct estimates that vary within the Fréchet bounds, as in Figure 1, by harnessing the parametric structure that follows directly from the randomization process. Parametric structure within randomized experiments is easy to justify as long as it corresponds to the actual randomization process, which is under the control of the people who design and implement the experiment. In many applications, economists use the parameteric structure of a model to infer aggregate counterfactuals from observed individual-level heterogeneity. We invert this procedure: using aggregate counterfactuals observed from the experiment, we apply the parametric structure of the randomization process to infer individual-level heterogeneity.
Our use of the likelihood implied by the design-based model is central to our contributions. The design-based model defines counterfactuals using causal models of potential outcomes from neyman1923, welch1937, kempthorne1952, copas1973, rubin1974; rubin1977, greenland1986, holland1986, and others, and then uses the randomization structure to derive the design-based likelihood. While copas1973 derives the design-based likelihood and uses it to test hypotheses about the average effect and the Fisher null hypothesis, our advance is to use the likelihood to learn about a more primitive object: the joint distribution of potential outcomes in the sample, including the number of defiers. We revisit Copas’ design-based model in light of modern causal models, modern assumptions about the absence of defiers, and modern computational advances. Likelihood-based models can provide exact distributions of experimental data, test statistics, and confidence intervals, but they remain uncommon relative to simulation methods in randomization inference. One exception is ding2019, who use the design-based likelihood to examine the sensitivity of inference on aggregate parameters to assumptions about the joint distribution of potential outcomes. rigdon2015 and li2016 construct exact confidence intervals without the design-based likelihood.
Our use of a decision-theoretic rule in the style of wald1949, rather than a classical hypothesis test, facilitates our progress. Classical hypothesis tests specify a null hypothesis and place the burden of evidence on the alternative hypothesis. This asymmetric treatment of Type I and Type II errors can imply an unrealistically conservative decision maker (tetenov2012) and exclude statistical evidence that, while weak, is still informative. To make use of all the evidence in our design-based likelihood, we focus on the statistical decision problem of choosing the correct joint distribution of potential outcomes in the sample, rather than on testing hypotheses about a prespecified distribution under prespecified size and power. We do not claim to provide a “test” of a monotonicity assumption. We provide a “decision” about the joint distribution of potential outcomes that can provide evidence for or against a monotonicity assumption.
Our use of design-based statistical decision theory contributes to the integration of statistical decision theory into econometrics (manski2004; dehejia2005; manski2007; hirano2008; hirano2009; stoye2012; kitagawa2018; manski2018; manski2019; manski2021; fernandez2024). Econometrics is a natural discipline to embrace statistical decision theory because statistical decision theory has such a tight relationship with economic theory. Statistical decision theory is natural to apply in our design-based setting because it does not require large sample assumptions or asymptotic approximations. Our use of design-based statistical decision theory departs from “asymptotic analysis of statistical decision rules in econometrics” as reviewed in hirano2020 in favor of the literature that focuses on statistical decision theory in finite sample settings (canner1970; manski2007admissible; schlag2007; stoye2007; stoye2009; tetenov2012).
We justify our proposed decision rule with justifications commonly employed in statistical decision theory, starting with the justification that our decision varies systematically with the data, as we show in Figure 3. One challenge with statistical decision criteria such as minimax and minimax regret is that researchers have shown instances in which they yield rules that do not vary with the data (schlag2003; manski2004; hirano2009; stoye2009). In his influential introduction to statistical decision theory, ferguson1967 recognizes that because maximum likelihood rules vary with the data, they are “reasonable” because they are better than “just guessing.” “However,” he continues, “decision theory, as developed here, is devoted to the problem of finding optimal rules, so that we do not refer to maximum likelihood estimates […] unless they turn out naturally to be optimal in some sense.”
In Section 2.4.2, we further justify our decision rule as Bayes optimal under the appropriate conditions, which advances Bayesian decision theory (chamberlain2011) by incorporating the design-based likelihood for our class of experiments. Bayes optimality implies that our rule is admissible—that is, there is no alternative rule that performs at least as well for all possible joint distributions of potential outcomes, and strictly better for some (ferguson1967). Furthermore, it ensures that our rule cannot be bested in a betting framework (freedman1969). The Bayesian interpretation of our rule facilitates our auxiliary construction of credible sets on the joint distribution of potential outcomes in Table E.1, providing an exact Bayesian alternative to classical asymptotic inference on Fréchet bounds (manski1992; horowitz2000; tamer2004; jun2023) and the number of defiers within the Fréchet bounds (imbens2004).
In addition to standard decision theoretic justifications, we also offer a perspective that our design-based maximum likelihood decision rule is optimal through the principle of maximum entropy (jaynes1957a; jaynes1957b), which does not require specification of a prior. golan2002 reviews a literature in econometrics that uses the principle of maximum entropy to abstract away from functional form assumptions on the likelihood function. In our setting, the design of the randomization, which is under the control of the experimenter, determines the functional form of the likelihood function, which also determines the functional form of the entropy.
Our results provide evidence that can support a monotonicity assumption or a specific alternative, which can be useful because monotonicity assumptions have become widespread, but they are not always realistic. Influential work by imbens1994 proposes a monotonicity assumption that assumes away compliers or defiers in the first stage of an instrumental variable model, and manski1997monotone proposes an analogous assumption in the reduced form. Monotonicity assumptions are ubiquitous in the analysis of first stages, and they are gaining traction in the analysis of reduced forms (alsan2025). First stage monotonicity assumptions are useful for the interpretation of instrumental variable estimates as local average treatment effects on compliers, and we view them as reasonable in many contexts. However, they can be difficult to defend in some contexts, particularly in health care settings that motivate our work because interventions may be helpful for some but harmful for others. We interpret our binary outcome as “takeup” in our stylized example and applications so that we can use the familiar terminology of always and never “takers” developed for the first stage (angrist1996), but the design-based likelihood is valid for other binary outcomes, like survival. Given concerns about side effects in medicine, we see potential for useful applications that have perfect compliance and a single binary outcome that represents survival. In those applications, evidence of defiers could motivate researchers to reduce the dosage of a potentially toxic intervention.
Previous approaches have revealed evidence that violates monotonicity assumptions, but only with additional data beyond our binary intervention and outcome. Machine learning approaches of wager2018 and semenova2024 require data on covariates to reveal heterogeneous effects. Analysis of side effects in medicine requires data on secondary outcomes (e.g. bernard2001). Specification tests of instrumental variable model assumptions can reveal evidence against monotonicity in the first stage (imbens1997; richardson2010; huber2012; huber2015testing; kitagawa2015; mourifie2017; machado2019), and marginal treatment effect methods can reveal evidence against monotonicity in the second stage (bjorklund1987; heckman1999; kowalski2023behavior; kowalski2023reconciling), but such approaches require the additional structure of a two-stage model as well as data on both stages. Evidence against first stage monotonicity by chan2022 requires a multi-valued instrument and the ability to observe a second stage outcome for one value of the first stage outcome. We contribute to the literature on violations of monotonicity by demonstrating that, even without additional data, our design-based model of an experiment can reveal evidence for or against monotonicity.
We view the development of our visualizations as a secondary contribution. Figure 1 allows us to depict how the same data can arise from different distributions of potential outcomes within a Fréchet set. Our visualization of how the MLE varies with the data in Figure 3 allows us to depict all possible results from an experiment of a given size, demonstrating that we have not engaged in selective reporting. The visualization allows us to characterize patterns that would be difficult to infer from isolated empirical applications and demonstrates the value of statistical decision theory over testing.
We also offer substantive insights in real empirical applications. In one, a payment intervention to induce pregnant women to quit smoking, we find no defiers. In another, a nudge intervention to get more people to hypothetically donate their organs, we find both compliers and the upper Fréchet bounds on defiers. In both applications, the weak evidence that we find can motivate stronger assumptions that researchers can use to learn more.
5 Implications for Econometric and Applied Research
We see our demonstration that a design-based model can reveal evidence beyond the average effect with very minimal assumptions and data as a proof-of-concept that could pave the way for the production of stronger evidence that will achieve the ultimate goal of personalizing decisions. Even if the average effect is the decision-relevant parameter in some contexts, learning about heterogeneity that underlies the average effect is a first step toward targeting that can ultimately yield a larger average effect. In our second empirical application, under the maximum likelihood joint distribution of potential outcomes, which includes no never takers, we can deduce that the people who are untreated in intervention must be defiers, and the people who are untreated in control must be compliers. Under monotonicity, it is never possible to deduce that a given individual is a complier, and defiers are assumed away. Our hope is that the ability to look directly at some compliers and defiers under alternative assumptions motivated by our statistical decision rule will pave the way to characterize them, target interventions toward compliers and away from noncompliers, and improve the average effect of future interventions.
Our work suggests several areas for further econometric research. Design-based likelihoods can be derived for more complex experimental settings, including those with two stages, multi-valued variables, and alternative randomization designs, such as matched pair, stratified, randomized permuted block, and designs that allow for rerandomization. A promising direction is applying design-based likelihoods to parameters traditionally studied under the lens of partial identification, such as the average treatment effect in experiments with imperfect compliance, missing data, attrition, spillovers, and measurement error. Another is using design-based likelihoods and entropy to motivate alternative parametric assumptions in auctions (jun2024) and other designed markets (abdulkadiroğlu2017). Approximation methods and advances in algorithmic efficiency could facilitate calculations in a wide array of samples. Optimal estimators under alternative priors and utility functions could be implemented with straightforward modifications. We are particularly interested in extending decision problems with “do no harm” objectives (mueller2023; benmichael2024; christy2024a; koch2025; gelman2025) to incorporate evidence from the design-based likelihood. Estimators could also be developed for alternative estimands, like potential outcome quantiles (cui2023; guggenberger2024). Finally, while previous results on optimal experimental design focus on estimating the average effect (bai2022), other designs, especially those that incorporate additional data, could reveal stronger evidence on the joint distribution of potential outcomes.
Design-based methods have been around for a long time, and we have demonstrated that it is feasible to implement our design-based decision rule, but a practical impediment to widespread implementation is that applied researchers often do not report enough information about their experiments to replicate them with design-based methods (young2019). We use large language models to pull (openAI’s GPT-4o-mini) and categorize (openAI’s GPT-o3-mini) the randomization processes from 2,080 papers associated with randomized controlled trials from the Abdul Latif Jameel Poverty Action Lab (J-PAL). Only 61% have enough description for the large language model to confidently categorize the randomization method from the paper or associated AEA RCT Registry entry. However, of those, 78% use designs captured by design-based likelihoods that we consider. In the full set of papers, over 60% describe some mechanism through which there could be defiers. Another practical impediment to design-based methods is that instead of reporting absolute numbers, several researchers only report rates or regression results. Mandated reporting of experimental design and standard sample counts by J-PAL, medical journals, and other authorities could facilitate the application of our design-based decision rule, allowing researchers to learn more from their experiments.
Appendix
Appendix A  Derivation of the Design-Based Likelihood
The design-based likelihood of the joint distribution of potential outcomes in the sample ${\boldsymbol{\theta}}=(\theta_{11},\theta_{10},\theta_{01},\theta_{00})$ given the data $\boldsymbol{x}$} is the probability of this realization of the data given $\boldsymbol{\theta}$}. First, we define the joint distribution of potential outcomes in the intervention arm ${\boldsymbol{I}}\equiv\big(I_{11},I_{10},I_{01},I_{00}\big)$ as a random vector whose elements represent the numbers of always takers, compliers, defiers, and never takers randomized into intervention; we write its realizations as ${\boldsymbol{i}}=(i_{00},i_{10},i_{01},i_{00})$. These values are unobserved to the experimenter. However, the randomization process—the design of the experiment—implies a data generating process for the distribution of potential outcomes in the intervention arm $\boldsymbol{I}$} conditional on the distribution of potential outcomes in the sample $\boldsymbol{\theta}$}, which we can then use to derive the distribution of the data $\boldsymbol{X}$}.
Consider an experiment where the randomization design assigns subjects to the intervention or control arm through a series of Bernoulli trials. Then, each subject’s probability of assignment to intervention is $p$, and assignments are independent across subjects as well as across principal strata. This independence allows us to write the distribution of $\boldsymbol{I}$ as the product of four independent binomial distributions:
${\mathbb{P}}\Big({\boldsymbol{I}}={\boldsymbol{i}}\mid{\boldsymbol{\theta}}\Big)=\binom{\theta_{11}}{i_{11}}\binom{\theta_{10}}{i_{10}}\binom{\theta_{01}}{i_{01}}\binom{\theta_{00}}{i_{00}}p^{\sum_{j,k}i_{j,k}}(1-p)^{n-\sum_{j,k}i_{j,k}}.$ (5)
Alternatively, in a completely randomized experiment, the randomization design fixes the number of subjects in the intervention arm $m$ and selects any of the possible combinations of $m$ subjects in intervention and nmn-m subjects in control with equal probability, as though drawing names from a hat. Under this design, $\boldsymbol{I}$ follows a multivariate hypergeometric distribution:
${\mathbb{P}}\Big({\boldsymbol{I}}={\boldsymbol{i}}\mid{\boldsymbol{\theta}}\Big)=\frac{\binom{\theta_{11}}{i_{11}}\binom{\theta_{10}}{i_{10}}\binom{\theta_{01}}{i_{01}}\binom{\theta_{00}}{i_{00}}}{\binom{n}{m}}$. (6)
While the distribution of potential outcomes in the intervention arm $\boldsymbol{I}$} is unobservable, we can use its data generating process to derive the distribution of the observable data $\boldsymbol{X}$}. Note that each subject randomized into the intervention arm with outcome $Y=1$ must be either an always taker or a complier: $X_{I1}=I_{11}+I_{10}$. Each subject randomized into the intervention arm with outcome $Y=0$ must be either a never taker or a defier: $X_{I0}=I_{00}+I_{01}$. In the control arm, those observed with outcome $Y=1$ must be either always takers that were not randomized into intervention, or defiers that were not randomized into intervention: $X_{C1}=\theta_{11}-I_{11}+\theta_{01}-I_{01}$. And finally, those in the control arm with outcome $Y=0$ must be either never takers that were not randomized into intervention, or compliers that were not randomized into intervention: $X_{C0}=\theta_{00}-I_{00}+\theta_{10}-I_{10}$. Thus, we can write the probability of the observed data $\boldsymbol{X}$} conditional on the joint distribution of potential outcomes $\boldsymbol{\theta}$} as:
${\mathbb{P}}\big({\boldsymbol{X}}={\boldsymbol{x}}\mid{\boldsymbol{\theta}}\big)$ $={\mathbb{P}}\Big(I_{11}+I_{10}=x_{I1},\ \big(\theta_{11}-I_{11}\big)+\big(\theta_{01}-I_{01}\big)=x_{C1},$
$\qquad\qquad I_{00}+I_{01}=x_{I0},\ \big(\theta_{00}-I_{00}\big)+\big(\theta_{10}-I_{10}\big)=x_{C0}\mid{\boldsymbol{\theta}}\Big)$
$={\mathbb{P}}\Big(I_{11}+I_{10}=x_{I1},\ I_{11}+I_{01}=x_{C1}-\theta_{11}-\theta_{01},$
$\qquad\qquad I_{00}+I_{01}=x_{I0},\ I_{00}+I_{10}=x_{C0}-\theta_{00}-\theta_{10}\mid{\boldsymbol{\theta}}\Big).$
A realization of the data $\boldsymbol{X}$} may be produced from more than one realization of the distribution of potential outcomes in intervention $\boldsymbol{I}$}. To find the probability of a realization of $\boldsymbol{X}$}, we sum together the probabilities of each realization of $\boldsymbol{I}$} that could have produced it. We can index these realizations through the realization $i$ of $I_{11}$ and solve the following system of equations for the elements of $\boldsymbol{I}$}:
$I_{11}+I_{10}$ $=x_{I1},$
$I_{11}+I_{01}$ $=\theta_{11}+\theta_{01}-x_{C1},$
$I_{11}+I_{10}+I_{01}+I_{00}$ $=x_{I1}+x_{I0},$
$I_{11}$ $=i.$
Rearranging yields
$I_{11}$ $=i$
$I_{10}$ $=x_{I1}-i$
$I_{01}$ $=\theta_{11}+\theta_{01}-x_{C1}-i$
$I_{00}$ $=x_{I0}+x_{C1}+i-\theta_{11}-\theta_{01}.$
The value $i$ is restricted to the set $\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})$ such that $\boldsymbol{I}$} remains within the support implied by $\boldsymbol{\theta}$}, namely $0\leq I_{11}\leq\theta_{11}$, $0\leq I_{10}\leq\theta_{10}$, $0\leq I_{01}\leq\theta_{01}$, and $0\leq I_{00}\leq\theta_{00}$. The probability of a realization of $\boldsymbol{X}$} is just the sum of the probability of each of these realizations of $\boldsymbol{I}$}:
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})={\mathbb{P}}\big({\boldsymbol{X}}={\boldsymbol{x}}\mid{\boldsymbol{\theta}}\big)=\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}{\mathbb{P}}\Big(I_{11}$ $=i,$
$I_{10}$ $=x_{I1}-i,$
$I_{01}$ $=\theta_{11}+\theta_{01}-x_{C1}-i,$
$I_{00}$ $=x_{I0}+x_{C1}+i-\theta_{11}-\theta_{01}\mid{\boldsymbol{\theta}}\Big).$ (7)
Equation (7) represents the general design-based likelihood, where the probability of each distribution of potential outcomes in intervention $\boldsymbol{I}$} given the distribution of potential outcomes in the sample $\boldsymbol{\theta}$} is informed by the design of the experiment. When subjects are assigned to intervention by a series of Bernoulli trials, $\boldsymbol{I}$} follows the distribution in (5), yielding:
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $=\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{i}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-i}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-i}$
$\qquad\qquad\times\binom{\theta_{00}}{x_{I0}+x_{C1}+i-\theta_{11}-\theta_{01}}$
$\qquad\qquad\times p^{x_{I1}+x_{I0}}(1-p)^{x_{C1}+x_{C0}}.$ (8)
Alternatively, in a completely randomized experiment, $\boldsymbol{I}$} follows the distribution in (6), yielding:
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $=\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{i}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-i}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-i}$
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}\bigg/\binom{n}{m}$. (9)
where we have substituted $m=x_{I1}+x_{I0}$. In either case, once conditioning on the realized data $\boldsymbol{x}$}, the likelihood is proportional to
$\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})$ $\propto\sum_{i\in\mathcal{I}({\boldsymbol{x}},{\boldsymbol{\theta}})}\binom{\theta_{11}}{i}$
$\qquad\qquad\times\binom{\theta_{10}}{x_{I1}-i}$
$\qquad\qquad\times\binom{\theta_{01}}{\theta_{11}+\theta_{01}-x_{C1}-i}$
$\qquad\qquad\times\binom{\theta_{00}}{m+x_{C1}+i-\theta_{11}-\theta_{01}-x_{I1}}$. (10)
This final expression is equivalent to the likelihood in copas1973.
Appendix B  Derivation of a Canonical Sampling-Based Likelihood
Suppose that subjects are sampled into the experiment with replacement from a population: let $q_{11}$ represent the share of the population who are always takers, $q_{10}$ the share of compliers, $q_{01}$ the share of defiers, and $q_{00}$ the share never takers. Sampling with replacement implies that each subject’s pair of potential outcomes is independently and identically distributed (IID) according to these shares.
We can derive likelihood functions for the population distribution of potential outcomes ${\boldsymbol{q}}\equiv(q_{11},q_{10},q_{01},q_{00})$ given the data from the experiment, depending on the randomization procedure into intervention. If subjects are assigned to intervention or control via a series of Bernoulli trials, we can view the combination of assignment $Z$ and outcome $Y$ across the $n$ subjects as independent categorical variables whose distribution is defined by the population distribution of potential outcomes $\boldsymbol{q}$} and the probability of assignment to intervention $p$:
${\mathbb{P}}\big(Y=1,Z=I\big)$ $=(q_{11}+q_{10})p$
${\mathbb{P}}\big(Y=0,Z=I\big)$ $=\big(1-(q_{11}+q_{10})\big)p$
${\mathbb{P}}\big(Y=1,Z=C\big)$ $=(q_{11}+q_{01})(1-p)$
${\mathbb{P}}\big(Y=0,Z=C\big)$ $=\big(1-(q_{11}+q_{01})\big)(1-p).$
The experimental data $\boldsymbol{X}$} are counts of realizations of independent categorical variables. Therefore, $\boldsymbol{X}$} follows the multinomial distribution with the probabilities above, which yields the likelihood expression:
$\mathcal{L}({\boldsymbol{q}}\mid{\boldsymbol{x}})$ $={\mathbb{P}}({\boldsymbol{X}}={\boldsymbol{x}}\mid{\boldsymbol{q}})$
$=\frac{n!}{x_{I1}!x_{I0}!x_{C1}!x_{C0}!}p^{x_{I1}+x_{I0}}(1-p)^{x_{C1}+x_{C0}}$
$\qquad\times\big(q_{11}+q_{10}\big)^{x_{I1}}\big(1-(q_{11}+q_{10})\big)^{x_{I0}}$
$\qquad\times\big(q_{11}+q_{01}\big)^{x_{C1}}\big(1-(q_{11}+q_{01})\big)^{x_{C0}}.$ (11)
This likelihood is equivalent to a likelihood appearing in barnard1947.
Alternatively, if the experiment is “completely randomized,” such that $m\leq n$ subjects are assigned to intervention by drawing names from a hat and the remainder are assigned to control, we can view the sample as two sets of Bernoulli trials. The first set of $m$ trials represents the subjects sampled into the intervention arm, where there are $X_{I1}$ treated subjects with independent probabilities of being treated equal to the probability that $Y_{I}=1$ in the population, $q_{11}+q_{10}$. The second set of nmn-m trials represents the subjects sampled into the control arm, where there are $X_{C1}$ treated subjects with independent probabilities of being treated equal to the probability that $Y_{C}=1$ in the population, $q_{11}+q_{01}$. The two sets of Bernoulli trials are independent, allowing us to write the likelihood of a completely randomized experiment as the product of two Binomial random variables $X_{I1}$ and $X_{C1}$:
$\mathcal{L}({\boldsymbol{q}}\mid{\boldsymbol{x}})$ $=\binom{m}{x_{I1}}\big(q_{11}+q_{10}\big)^{x_{I1}}\big(1-(q_{11}+q_{10})\big)^{m-x_{I1}}$
$\qquad\times\binom{n-m}{x_{C1}}\big(q_{11}+q_{01}\big)^{x_{C1}}\big(1-(q_{11}+q_{01})\big)^{n-m-x_{C1}}.$ (12)
This likelihood is equivalent to likelihoods from barnard1947 and kline2020.
In each of these likelihood functions, the joint distribution of potential outcomes in the population $\boldsymbol{q}$} appears only through the transformations $q_{11}+q_{10}$ and $q_{11}+q_{01}$, which are the marginal distributions of the potential outcomes in intervention and control in the population, respectively. Therefore, the likelihoods of any two distributions in the same population-level Fréchet set are equal. That is, the experimental data contains no information about the joint distribution of potential outcomes in the population beyond the marginal distributions of potential outcomes.
Appendix C  Bayes Optimality of Our Proposed Rule
Appendix C.1  Preliminaries on Statistical Decision Theory
Suppose a decision maker wishes to guess the joint distribution of potential outcomes in the sample. We write the decision maker’s guess as $\widehat{{\boldsymbol{\theta}}}$, and we define a utility function over a guess $\widehat{{\boldsymbol{\theta}}}$ and the true distribution $\boldsymbol{\theta}$}. We focus on the following utility specification, which yields one util when the guess is correct and zero utils when the guess is incorrect:
$u(\widehat{{\boldsymbol{\theta}}},{\boldsymbol{\theta}})=\mathbf{1}_{\{\widehat{{\boldsymbol{\theta}}}={\boldsymbol{\theta}}\}}.$
We also allow the decision maker to choose a randomized guess, which ascribes a probability distribution over the possible values of $\boldsymbol{\theta}$}. We define the decision maker’s utility over a randomized guess $\rho$ as the expected utility of guessing according to the probabilities ascribed by $\rho$:
[…]
The decision maker chooses a decision rule that maps the observable data into (possibly) randomized guesses. [^3] We conflate here the standard definitions of “randomized decision rules” and “behavioral decision rules” (ferguson1967) for expositional clarity. In settings of perfect recall, such as the setting we study here, the space of randomized and behavioral decision rules is equivalent (kuhn1953). We write such a rule as $f:\boldsymbol{\mathcal{X}}\to\Delta({\boldsymbol{\Theta}})$, where $\boldsymbol{\mathcal{X}}$ is the space of possible data realizations, $\boldsymbol{\Theta}$} is the space of possible distributions of potential outcomes, and $\Delta({\boldsymbol{\Theta}})$ is the space of distributions over $\boldsymbol{\Theta}$}. Given a true distribution of potential outcomes $\boldsymbol{\theta}$}, the decision maker’s expected utility from following a decision rule $f$ is the expected value of $U(f({\boldsymbol{X}}),{\boldsymbol{\theta}})$ with respect to the experimental outcome $\boldsymbol{X}$}:
$EU(f,{\boldsymbol{\theta}})$ $=\mathbb{E}\big[U\big(f({\boldsymbol{X}}),{\boldsymbol{\theta}}\big)\mid{\boldsymbol{\theta}}\big]$
[…]
$=\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})f({\boldsymbol{x}})({\boldsymbol{\theta}}).$
The last line follows from our chosen utility function, under which the decision maker’s expected utility is equal to the probability of guessing correctly.
One sense in which a decision rule can be optimal is in terms of Bayes expected utility: the average expected utility obtained according to some subjective prior distribution over $\boldsymbol{\theta}$}:
$EU_{\pi}(f)=\mathbb{E}_{\pi}\big[EU(f,{\boldsymbol{\theta}})\big]=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}EU(f,{\boldsymbol{\theta}})\pi({\boldsymbol{\theta}}),$ (14)
where $\pi\in\Delta({\boldsymbol{\Theta}})$ is a subjective prior belief about $\boldsymbol{\theta}$}. Decision rules that maximize this criterion are said to be “Bayes optimal.”
Appendix C.2  Conditions for Bayes Optimality
In this appendix, we review the result that the maximum likelihood decision rule is Bayes optimal when utility takes the form in (13) and the decision maker’s subjective prior is uniform across realizations of the distribution of potential outcomes in the sample $\boldsymbol{\theta}$}. To show this result, we first establish that the maximum a posteriori decision rule is Bayes optimal for our chosen loss function given any prior. The maximum a posteriori decision rule $f^{*}_{\pi}$ selects the maxima of the posterior distribution of $\boldsymbol{\theta}$}. Formally, let $\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})$ be the set of $\boldsymbol{\theta}$} values that maximize the posterior distribution given the observed data $\boldsymbol{x}$}, i.e.
$\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})$ $={\underset{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}{\mathrm{arg}\,\mathrm{max}}\ }{\mathbb{P}}\big({\boldsymbol{\theta}}\mid{\boldsymbol{x}}\big)$
$={\underset{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}{\mathrm{arg}\,\mathrm{max}}\ }\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})\pi({\boldsymbol{\theta}}),$ (15)
where $\pi\in\Delta({\boldsymbol{\Theta}})$ is the prior belief about $\boldsymbol{\theta}$}. The decision rule $f^{*}_{\pi}$ can then be defined as follows:
$f^{*}_{\pi}({\boldsymbol{x}})({\boldsymbol{\theta}})=\begin{cases}\dfrac{1}{\#\{\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})\}}&\text{if}\ {\boldsymbol{\theta}}\in\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}}),\\ 0&\text{o.w.}\end{cases}$ (16)
Now, let $g$ be an arbitrary decision function. The Bayes expected utility for decision function $g$ is
${\mathbb{E}}\big[EU(g,{\boldsymbol{\theta}})\big]$ $=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}EU(g,{\boldsymbol{\theta}})\pi({\boldsymbol{\theta}})$
$=\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}\Bigg[\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})g({\boldsymbol{x}})({\boldsymbol{\theta}})\Bigg]\pi({\boldsymbol{\theta}}).$
By rearranging terms in the summation, we can bound the Bayes expected utility of $g$:
$\mathbb{E}\big[EU(g,{\boldsymbol{\theta}})\big]$ $=\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}\bigg(\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})g({\boldsymbol{x}})({\boldsymbol{\theta}})\pi({\boldsymbol{\theta}})\bigg)$
$\leq\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\ \sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}\bigg(g({\boldsymbol{x}})({\boldsymbol{\theta}})\max_{{\boldsymbol{\theta}}^{\prime}\in{\boldsymbol{\Theta}}}\Big\{\mathcal{L}({\boldsymbol{\theta}}^{\prime}\mid{\boldsymbol{x}})\pi({\boldsymbol{\theta}}^{\prime})\Big\}\bigg)$
$=\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\Bigg[\max_{{\boldsymbol{\theta}}^{\prime}\in{\boldsymbol{\Theta}}}\Big\{\mathcal{L}({\boldsymbol{\theta}}^{\prime}\mid{\boldsymbol{x}})\pi({\boldsymbol{\theta}}^{\prime})\Big\}\underbrace{\bigg(\sum_{{\boldsymbol{\theta}}\in{\boldsymbol{\Theta}}}g({\boldsymbol{x}})({\boldsymbol{\theta}})\bigg)}_{=1}\Bigg].$
This bound is precisely the Bayes expected utility achieved by decision rule $f^{*}_{\pi}$:
${\mathbb{E}}\big[EU(f^{*}_{\pi},{\boldsymbol{\theta}})\big]$ $=\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\sum_{{\boldsymbol{\theta}}\in\Theta}\bigg(\mathcal{L}({\boldsymbol{\theta}}\mid{\boldsymbol{x}})f^{*}_{\pi}({\boldsymbol{x}})({\boldsymbol{\theta}})\pi({\boldsymbol{\theta}})\bigg)$
$=\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\sum_{{\boldsymbol{\theta}}\in\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})}\bigg(\frac{1}{\#\{\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})\}}\max_{{\boldsymbol{\theta}}^{\prime}\in{\boldsymbol{\Theta}}}\Big\{\mathcal{L}({\boldsymbol{\theta}}^{\prime}\mid{\boldsymbol{x}})\pi({\boldsymbol{\theta}}^{\prime})\Big\}\bigg)$
$=\sum_{{\boldsymbol{x}}\in\boldsymbol{\mathcal{X}}}\Bigg[\max_{{\boldsymbol{\theta}}^{\prime}\in{\boldsymbol{\Theta}}}\Big\{\mathcal{L}({\boldsymbol{\theta}}^{\prime}\mid{\boldsymbol{x}})\pi({\boldsymbol{\theta}}^{\prime})\Big\}\underbrace{\bigg(\sum_{{\boldsymbol{\theta}}\in\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})}\frac{1}{\#\{\widehat{{\boldsymbol{\Theta}}}_{\pi}({\boldsymbol{x}})\}}\bigg)}_{=1}\Bigg].$
Thus, since $f^{*}_{\pi}$ achieves the upper bound on the Bayes expected utility of any decision rule, we conclude that $f^{*}_{\pi}$ is Bayes optimal. Finally, observe that when the prior distribution $\pi(\theta)$ is constant, the maximizers of the posterior distribution in (15) are simply the maximizers of the likelihood, and the Bayes rule $f^{*}_{\pi}$ in (16) simplifies to our maximum likelihood decision rule.
For alternative utility functions or priors, the optimal rule may not take a convenient form such as the MLE, but can still be computed by solving the linear programming problem of maximizing Bayes expected utility in (14) with respect to the (finitely-many) values of $f({\boldsymbol{x}})({\boldsymbol{\theta}})$, such that $\sum_{\boldsymbol{\theta}}f({\boldsymbol{x}})({\boldsymbol{\theta}})=1$ for each $\boldsymbol{x}$}.
Appendix D  Construction of Credible Sets
We construct credible sets on the joint distribution of potential outcomes in the sample using the posterior belief implied by a uniform prior and the design-based likelihood. A credible set of level $\alpha$ is a set of joint distributions of potential outcomes covering a posterior probability mass of at least $\alpha$. We report the smallest credible set, which is the credible set containing the fewest points. We construct these sets with the following algorithm:
  1. 1.
    Compute the posterior probability of every joint distribution of potential outcomes $\boldsymbol{\theta}$} given the observed data $\boldsymbol{x}$}.
  2. 2.
    Order the joint distributions from highest to lowest posterior probability.
  3. 3.
    Compute the cumulative probability mass according to this ordering.
  4. 4.
    Starting with the highest probability distribution (which is the MLE), add joint distributions of potential outcomes to the credible set until the probability mass of the credible set exceeds $\alpha$. In the case of ties, we add all distributions with the same probability simultaneously.
To report credible sets on the number of always takers, compliers, defiers, or never takers individually, we report the range of values for that type present in the credible set constructed on the joint distribution. Note that the credible sets on individual response types are not independent. For example, the maximal values in the individual credible sets for always takers, compliers, defiers, and never takers may not constitute an element of the credible set on the joint distribution of potential outcomes.
Appendix E  Auxiliary Statistics
Figure E.1: Standard Statistics, Proposed Design-Based Maximum Likelihood Estimates, and Auxiliary Statistics